If the perimiter of a rectangle is 78 units and the lenght is 5 units larger than the width. Whats is the length?
Answer by Cromlix(4381)

Width = x
Length = x + 5
Perimeter = 2*x + 2*(x + 5)
78 = 2x + 2x + 10
Collect like terms;
4x = 78 - 10
4x = 68
x = 17
Width = 17 units
Length = 22 units
Hope this helps.
:-)